### v1

# MSM_pseudo_CS

Abstract
Case-I interval-censored (current status) data from multistate systems are often encountered in cancer and other epidemiological studies. In this article, we focus on the problem of estimating state entry distribution and occupation probabilities, contingent on a preceding state occupation. This endeavor is particularly complex owing to the inherent challenge of the unavailability of directly observed counts of individuals at risk of transitioning from a state, due to the cross-sectional nature of the data. We propose two nonparametric approaches, one using the fractional at-risk set approach recently adopted in the right-censoring framework and the other a new estimator based on the ratio of marginal state occupation probabilities. Both estimation approaches utilize innovative applications of concepts from the competing risks paradigm. The finite-sample behavior of the proposed estimators is studied via extensive simulation studies where we show that the estimators based on severely censored current status data have good performance when compared with those based on complete data. We demonstrate the application of the two methods to analyze data from patients diagnosed with breast cancer.
1 Introduction
Multistate models are a popular type of survival model that allows for studying the progression of the disease using multivariate time-to-event data. In biomedical applications, situations arise in which patients progress reversibly or irreversibly through a succession of disease stages that describe their current condition. The patients usually begin from an initial disease state and move through one or more intermediate states before entering a final absorbing state. The disease stages correspond to the states of the multistate system. Moreover, in certain applications, the transition of patients through their multistate system may be subject to a severe form of (case-I) interval censoring often referred to as current status censoring (Jewell and van der Laan, 2003; Sun, 2006). Under such an observation scheme, the researcher is only restricted to information on the state occupied by the patient at a (single) random inspection time during the entire study period.
Current status data from multistate models often arise in epidemiological and biomedical research, as well as in observational studies where continuous monitoring of study participants is not feasible. With event history data from a multistate system, practitioners are often interested in estimating the probability of occupying a given state, at a specific time. Datta and Sundaram (2006) proposed marginal estimators for state occupation probabilities following the product integration approach of Aalen and Johansen (1978). Their nonparametric approach is valid without imposing the Markov assumption, and it has been further studied under different settings (Lan and Datta, 2010; Lan et al., 2017; Anyaso-Samuel et al., 2023).
In progressive multistate systems, where individuals navigate through successive states within a disease process, a pertinent question arises: “What is the probability of occupying a future state $k$, given a prior visit to state $j$?” We denote this conditional probability as $\Psi_{k|j}$. Furthermore, interest may extend to studying the probability of reaching a future state $k$ by time $t$, given a previous visit to state $j$, represented as $\Psi_{k|j}(t)$. Historically, these conditional probabilities have garnered less attention, as multistate systems are often assumed to adhere to Markovity. However, practitioners increasingly recognize the limitations of imposing a Markov assumption, which may be unverifiable and violated. Given a progressive multistate system, understanding the proportion of individuals on a unique path capable of occupying a future state becomes crucial. Our paper addresses this gap by focusing on the conditional estimation of future state occupation probabilities given prior occupation of a specific state, within the context of current status censoring.
In the context of right-censored data, Yang et al. (2023) recently proposed an estimator for the future state occupational probability conditional on a past state visit. Constructing such estimators is difficult since information on the transitions out of the non-absorbing states within the multistate system is only partially observed. With current status data, the problem becomes even more challenging since the path that individuals would have taken after they were observed is unknown. To tackle this problem of estimating $\Psi_{k|j}$ and $\Psi_{k|j}(t)$, we propose two estimation methods. The first is based on fractional at-risk sets. Here, we formulate the at-risk processes with weights defined by the fractional observation that a person contributes to the at-risk set of transitioning out of state $j$. We estimate this fractional weight as the probability that a patient will eventually reach state $j$ given that they were occupying a known state at a random inspection time. The idea for utilizing fractional weighted at-risk sets follows from the seminal work of Datta and Satten (2000) and the work of Yang et al. (2023). We begin the construction of the second estimator by providing insight that the conditional quantity $\Psi_{k|j}$ can be expressed as the ratio of two marginal state occupation probabilities for multistate models with a tree structure. This leads to the formulation of an estimator based on product limit estimators of state occupation probabilities (Datta and Sundaram, 2006). For both estimation methods, we do not impose the assumption that the multistate systems are Markov, but the methods require that the systems have a directed tree structure. Following our construction of estimators for the conditional future state occupation probabilities, we derive estimators for the distribution function of the state entry conditional on a prior occupation of a given state.
The rest of the paper is organized as follows. In Section 2, we introduce the notation along with marginal nonparametric estimators for current status multistate data. In Section 3, we discuss the estimation of $\Psi_{k|j}$ based on the fractional weight approach. The alternate estimator based on the product-limit method is discussed in Section 4. In Section 5, we describe the estimation of pointwise confidence intervals for the estimators based on both approaches. Detailed simulation studies to evaluate the performance of the proposed estimators are presented in Section 6. In Section 7, we illustrate the application of both estimation methods in analyzing data from a breast cancer study. The paper concludes with a discussion in Section 8.
2 Nonparametric estimators for a multistate system
2.1 Notation and Convention
Consider a multistate system with a directed tracking structure, where $\mathcal{S}=\{0,1,\ldots,Q\}$ represents the finite state space. In this system, each individual $i=1,\ldots,n$ occupies a state denoted by $S_{i}(t)\in\mathcal{S}$ at time $t\geq 0$. Additionally, $C_{i}$ stands for the random inspection time for individual $i$. For an individual, the observed current-status data is represented as the pair $\{C_{i},S_{i}(C_{i})\}$, where $C_{i}$ is assumed to be independent of the multistate process $\{S_{i}(t),t\geq 0\}$, based on the random censoring assumption. To model the multistate current-status data, we adopt a framework where each state $j\in\mathcal{S}$ can be reached from the initial state 00 through a unique path: $0=s_{1}\rightarrow s_{2}\rightarrow\ldots\rightarrow s_{(j)+1}=j$, where $(j)$ denotes the number of steps it takes to reach $j$ from the root node 0.
Let $U_{jj^{\prime}}$ denote the (unobserved) transition time at which an individual transitions from $j$ to $j^{\prime}$. To estimate temporal functions of the multistate system, two essential functionals are the counting process denoted by $N_{jj^{\prime}}(t)$ and the at-risk process denoted by $Y_{j}(t)$. The counting process $N_{jj^{\prime}}(t)$ counts the number of transitions from state $j$ to $j^{\prime}$ in the time interval $[0,t)$. On the other hand, the at-risk process $Y_{j}(t)$ represents the number of individuals who are at risk of transitioning out of state $j$ at time $t$. These processes are expressed as follows
$N_{jj^{\prime}}(t)$ $=\sum^{n}_{i=1}\mathcal{I}(U_{i,jj^{\prime}}\leq t),\quad\text{ and }$
$Y_{j}(t)$ $=\sum^{n}_{i=1}\mathcal{I}(S_{i}(t-)=j),$
where $\mathcal{I}(\cdot)$ denotes the indicator function and $X(t-)$ is the state occupied just before time $t$. By the law of large numbers,
[…]
for any $t\geq 0$. Since the inspection time $C$ is independent of the multistate process, we can write
$\text{Pr}(U_{jj^{\prime}}\leq t)=\mathbb{E}(\mathcal{I}(U_{jj^{\prime}}\leq C)\ |\ C=t),$
where $\mathcal{I}(U_{jj^{\prime}}\leq C)$ is the indicator of the event that the transition from state $j$ to $j^{\prime}$ has taken place by time $C$.
Using the theory of nonparametric regression, Datta and Sundaram (2006) proposed nonparametric estimators of $N_{jj^{\prime}}(t)$ and $Y_{j}(t)$ under this context of current status data. For $N_{jj^{\prime}}(t)$, they proposed a two-step estimation approach where first, isotonic regression via the pooled adjacent violators (PAV) algorithm (Barlow et al., 1972) is performed based on the pairs $(C_{i},\ \mathcal{I}(U_{i,jj^{\prime}}\leq C_{i}))$. Let $\widehat{N}^{P}_{jj^{\prime}}(t)$ denote the fitted responses from the isotonic regression. Then secondly, kernel smoothing (Mukerjee, 1988; Nadaraya, 1964; Watson, 1964) with a log-concave density, $K>0$ is performed to remove long flat parts of $\widehat{N}^{P}_{jj^{\prime}}(t)$. The resulting estimator of $N_{jj^{\prime}}(t)$ is given by
$\widehat{N}_{jj^{\prime}}(t)=\sum^{n}_{i=1}\left\{\widehat{N}^{P}_{jj^{\prime}}(t)\times\frac{K_{h}(C_{i}-t)}{\sum^{n}_{i=1}K_{h}(C_{i}-t)}\right\},$ (1)
where $K_{h}(\cdot)=h^{-1}K(\cdot/h)$ with a bandwidth sequence $0<h=h(n)\downarrow 0$ chosen by the criteria specified by Wand and Jones (1995). Moreover, to estimate $\text{Pr}(X(t-)=j)$, which represents the probability limit of $n^{-1}Y_{j}(t)$, Datta and Sundaram (2006) employed the kernel smoothing approach. Accordingly, the estimation of $Y_{j}(t)$ is given by:
$\widehat{Y}_{j}(t)=\sum^{n}_{i=1}\left\{\mathcal{I}(S_{i}(C_{i})=j)\times\frac{K_{h}(C_{i}-t)}{n^{-1}\sum^{n}_{i=1}K_{h}(C_{i}-t)}\right\},$ (2)
where it is assumed that $S_{i}(C_{i}-)=S_{i}(C_{i})$ with probability 1. The estimators outlined in (1) and (2) pave the way for feasible inference in multistate models with current status data. For example, Datta and Sundaram (2006) elaborates on the estimation of key marginal functionals, such as the transition probabilities and state occupation probabilities, employing the product limit approach described by Andersen et al. (1993).
This paper primarily proposes estimators for future state occupation probabilities and state entry distributions conditioned on a prior state visit. We employ the methodology described above to estimate the counting process and at-risk set, which serves as the foundational framework for constructing our proposed conditional estimators.
3 Estimators based on fractional at-risk sets
3.1 An illustration using a five-state model
To provide context for the aim of this study, let’s examine the progressive five-state illness-death model depicted in Figure 1. Our focus centers on estimating the probability of ever visiting state 3, given a prior stay in state 1 over an individual’s lifespan. We denote this conditional future occupation probability as $\Psi_{3|1}$. Estimating $\Psi_{3|1}$ poses a challenge due to incomplete observation of transition times. With current status data, we only have access to information about the state a patient occupies at a random inspection time.
0. Disease-free 1. Ill 2. Death without illness 3. Death after illness 4. Death from other causes State 0*
Figure 1: A five-state illness-death model.
Let $X_{ij}$ denote the (unobserved) indicator of the event that the $i$th individual would ever enter state $j$. In the typical current status framework, $X_{ij}$ will likely be unknown, making estimating the future entry states difficult. For instance, if the $i$th individual was observed to have $S_{i}(C_{i})=0$, then $X_{i1}$ will be unobserved. Suppose the $i$th individual transitioned to state 1, they will be eligible for the competing risks (state 3 and state 4), but we would not know this because $X_{i1}$ is not observed. To overcome the difficulty in estimating $\Psi_{3|1}$, first, we define a new state $0^{*}$ which is constructed by combining states 0, 1, and 2 (see Figure 1). Note that not everyone (i.e., those who transitioned to state 2) in state $0^{*}$ will be subject to the competing risks.
Next, we compute the fractional observation, denoted as $\phi_{i1}(C_{i},S_{i}(C_{i}))$, representing the contribution of the $i$th individual to the “at-risk” set for transitioning out of state 1. This fractional observation estimates the probability of the $i$th individual eventually reaching state 1 from state 0. Computing $\phi_{i1}(C_{i},S_{i}(C_{i}))$ is straightforward when $S_{i}(C_{i})\neq 0$. For $S_{i}(C_{i})=0$, we compute $\phi_{i1}(C_{i},S_{i}(C_{i}))$ as the transition probability of moving from state 0 to state 1, that is, $\phi_{i1}(C_{i},S_{i}(C_{i}))={P}_{01}(C_{i},\infty)$. Using the Aalen-Johansen formula for transition probability, we estimate $\phi_{i1}(C_{i},S_{i}(C_{i}))$ by the following expression
[…]
where $\widehat{N}_{\ell\ell^{\prime}}(t)$ and $\widehat{Y}_{\ell}(t)$ are obtained by (1) and (2), respectively. The calculation of (3) does not require the Markov assumption since it pertains to the transition out of the root node. If $S_{i}(C_{i})=2$, then $\widehat{\phi}_{i1}(C_{i},S_{i}(C_{i}))=0$. For $S_{i}(C_{i})\in\{1,3,4\}$, indicating that the individual followed the $0\rightarrow 1$ path, $\widehat{\phi}_{i1}(C_{i},S_{i}(C_{i}))=1$.
To obtain the estimator of $\Psi_{3|1}$, we introduce a modified multistate system that includes state 0* as the initial stage, leading to either state 3 or state 4 (see Figure 1). Applying the Aalen-Johansen formula, commonly used in competing risks calculations, to this modified multistate system yields an expression that serves as the estimator for eventually reaching state 3:
[…]
Here the superscript $[1]$ is used to distinguish this estimator from the alternate estimator discussed in Section 4. The estimators of the counting processes and the at-risk set in (4) are defined by
$\widehat{N}_{3|1}(t)$ $=\sum^{n}_{i=1}\left\{\frac{\widehat{N}^{P}_{13}(t)K_{h}(C_{i}-t)}{\sum^{n}_{i=1}K_{h}(C_{i}-t)}\right\},$
$\widehat{N}_{4|1}(t)$ $=\sum^{n}_{i=1}\left\{\frac{\widehat{N}^{P}_{14}(t)K_{h}(C_{i}-t)}{\sum^{n}_{i=1}K_{h}(C_{i}-t)}\right\},$
$\widehat{Y}_{0^{*}|1}(t)$ $=\sum^{n}_{i=1}\left\{\widehat{\phi}_{i1}(C_{i},S_{i}(C_{i}))\times\frac{\mathcal{I}(S_{i}(C_{i})\in\{0,1,2\})\times K_{h}(C_{i}-t)}{n^{-1}\sum^{n}_{i=1}K_{h}(C_{i}-t)}\right\},$
where $n^{-1}\widehat{N}^{P}_{jj^{\prime}}(\cdot)$ is a step function for taking values $n^{-1}\widehat{N}^{P}_{jj^{\prime}}(C_{(i)})=R_{i}$, that minimizes the sum of squares $\sum_{i=1}^{n}\Big{\{}R_{i}-\mathcal{I}(U_{[i],jj^{\prime}}\leq C_{(i)})\Big{\}}^{2}$ subject to $R_{1}\leq...\leq R_{n}$, and $[i]$ corresponds to the index such that $C_{[i]}$ equals the $i$th-order statistic $C_{(i)}$. Subsequently, to obtain, $\widehat{N}_{j^{\prime}|j}(\cdot)$, similar to the approach used in constructing (1), we use kernel smoothing to remove the long flat part from $\widehat{N}^{P}_{jj^{\prime}}(\cdot)$. Note, the estimator, $\widehat{Y}_{0^{*}|1}(t)$ is formulated by the estimate of the fractional weight, $\widehat{\phi}_{i1}(C_{i},S_{i}(C_{i}))$. Here, $\widehat{Y}_{0^{*}|1}(t)$ estimates the number of persons at risk of transitioning out of the artificial state $0^{*}$ at time $t$.
Additionally, the estimation of the entry time distribution into state 3, given a prior visit to state 1, is expressed as follows:
$\widehat{F}_{3|1}(t)=\frac{\widehat{\Psi}^{[1]}_{3|1}(t)}{\widehat{\Psi}^{[1]}_{3|1}}.$
Here, the estimation of the subdistribution function $\widehat{\Psi}^{[1]}_{3|1}(t)$ is achieved by constraining the integral’s bounds as defined in the expression provided in Equation 4. Therefore,
[…]
3.2 Extension to general multistate model with a tree structure
Next, we expand the explanation of the estimator for $\Psi_{k|j}=P(X_{k}=1~{}|~{}X_{j}=1)$ to encompass a broader multistate model context. The estimator (based on fractional weights) of the probability of occupying state $k$ given a prior visit to state $j$ is denoted as $\widehat{\Psi}^{[1]}_{k|j}$. The estimation of ${\Psi}_{k|j}$ relies on the assumption that state $j$ lies on the unique path connecting states 00 and $k$; otherwise, ${\Psi}_{k|j}=0$. Let $m$ represent the number of transitions needed to move from state 00 to state $k$. Constructing the estimator for $\Psi_{k|j}$ follows a recursive approach based on $m$.
When $m=1$, the computation of $\widehat{\Psi}^{[1]}_{k|j}$ involves using the Aalen-Johansen estimator for transition probabilities within a competing risks framework. On the other hand, when $m=2$, then, one can obtain $\widehat{\Psi}^{[1]}_{k|j}$ by following the procedure outlined in Section 3.1. More generally, for any state $k$ that is situated at a distance of $m$ from state 0, the estimation of ${\Psi}_{k|j}$ relies on the application of the chain rule of conditional probability, given that the states of interest are aligned along a distinct path originating from the root node.
Following the application of the chain rule, for every state $j$ on the path from state 0 to $k$, where $k$ is a distance $m$ from state 0; define
$\widehat{\Psi}^{[1]}_{k|j}=\widehat{\Psi}^{[1]}_{\tilde{k}|j}\cdot\widehat{\Psi}^{[1]}_{k|\tilde{k}}$ (6)
where $\tilde{k}$ is the state just before state $k$ on the path from 00 to $k$. From this recursive formula, $\widehat{\Psi}^{[1]}_{\tilde{k}|j}$ is available since the distance from state 0 to state $\tilde{k}$ is less than $m$. To construct the estimator $\widehat{\Psi}^{[1]}_{k|\tilde{k}}$, we utilize the idea of estimating transition probabilities in the competing risk framework and the concept of “pooling of states”. Thus, we have
$\widehat{\Psi}^{[1]}_{k|\tilde{k}}=\int^{\infty}_{0}\widehat{S}_{0^{*}|\tilde{k}}(s-)\frac{d\widehat{N}_{k|\tilde{k}}(s)}{\widehat{Y}_{0^{*}|\tilde{k}}(s)}.$ (7)
where
$\widehat{S}_{0^{*}|\tilde{k}}(t)$ $=\prod_{(0,t)}\left[1-\frac{d\widehat{N}_{.|\tilde{k}}(v)}{\widehat{Y}_{0^{*}|\tilde{k}}(v)}\right].$
Here, state $0^{*}$ denotes an artificial state constructed by combining the states preceding (and including) $\tilde{k}$. Also,
$\widehat{N}_{k|\tilde{k}}(t)$ $=\sum^{n}_{i=1}\left\{\frac{\widehat{N}^{P}_{\tilde{k}k}(t)K_{h}(C_{i}-t)}{\sum^{n}_{i=1}K_{h}(C_{i}-t)}\right\}\text{ and }\widehat{N}_{.|\tilde{k}}(t)=\sum_{k}\widehat{N}_{k|\tilde{k}}(t),$
$\widehat{Y}_{0^{*}|\tilde{k}}(t)$ $=\sum^{n}_{i=1}\left\{\widehat{\phi}_{i\tilde{k}}\times\frac{\mathcal{I}(S_{i}(C_{i})\in 0^{*})K_{h}(C_{i}-t)}{n^{-1}\sum^{n}_{i=1}K_{h}(C_{i}-t)}\right\},$
where $\widehat{N}^{P}_{\tilde{k}k}(\cdot)$ is obtained by the isotonic regression of the pairs $(C_{i},\ \mathcal{I}(U_{i,\tilde{k}k}\leq C_{i}))$. The estimator of the number of persons at risk in state $0^{*}$, $\widehat{Y}_{0^{*}|\tilde{k}}(t)$, is based on the fractional weight $\widehat{\phi}_{i\tilde{k}}(C_{i},S_{i}(C_{i}))$ that estimates the probability of individual $i$ reaching state $\tilde{k}$ based on the information observed at their inspection time $C_{i}$. We interpret $\widehat{\phi}_{i\tilde{k}}(C_{i},S_{i}(C_{i}))$ as the contribution of the $i$th individual to the “at-risk” set of transitioning out of state $\tilde{k}$. If individual $i$ is observed to occupy state $\tilde{k}$ at their inspection time, then, $\widehat{\phi}_{i\tilde{k}}(C_{i},S_{i}(C_{i}))=1$, otherwise if they were observed to occupy a state not on the path from state 0 to state $\tilde{k}$, then $\widehat{\phi}_{i\tilde{k}}(C_{i},S_{i}(C_{i}))=0$. The estimation of $\widehat{\phi}_{i\tilde{k}}(C_{i},S_{i}(C_{i}))=\widehat{\mathbb{P}}(X_{i\tilde{k}}=1\ |\ S_{i}(C_{i})=\ell)$ only becomes intricate when such individual occupies a state $\ell$ which lies on the path from state 0 to state $\tilde{k}$ at their inspection time. To calculate $\widehat{\phi}_{i\tilde{k}}(C_{i},S_{i}(C_{i}))$ for such cases, we utilize a competing risk calculation where the states prior of $\tilde{k}$ are combined to form an event-free state.
Further, we describe the estimation of the entry distribution for state $k$ given a previous occupancy of state $j$, denoted as $\widehat{F}_{k|j}(t)$. Let $\mathcal{S}^{k}$ denote the set of states $\ell$ such that state $k$ is on the path from state 0 to state $\ell$. Then, $\widehat{F}_{k|j}(t)$ can be expressed as:
$\widehat{F}_{k|j}(t)=\frac{\sum_{\ell\in\mathcal{S}^{k}}\widehat{\Psi}^{[1]}_{\ell|j}(t)}{\sum_{\ell\in\mathcal{S}^{k}}\widehat{\Psi}^{[1]}_{\ell|j}},$ (8)
where $\widehat{\Psi}^{[1]}_{k|j}(t)=\widehat{\Psi}^{[1]}_{\widetilde{k}|j}(t)\cdot\widehat{\Psi}^{[1]}_{k|\widetilde{k}}(t)$, as defined in Eqn. (6). Note that, $\widehat{F}_{k|j}(t)$ is the normalized sum of estimated future occupation probabilities (conditioned on a prior visit to state $j$) of state $k$ and all other states that come after state $k$ in the progressive system.
For a practical illustration, consider a seven-state multistate system (refer to Figure 4). Suppose we aim to estimate the probability of occupying state 5 given a prior visit to state 1, denoted by $\Psi_{5|1}$. Utilizing Eqn. (6), in this case, we have $k=5$, $j=1$, and $\widetilde{k}=3$. Consequently, our goal is to compute $\widehat{\Psi}^{[1]}_{5|1}=\widehat{\Psi}^{[1]}_{3|1}\cdot\widehat{\Psi}^{[1]}_{5|3}$. Following the recursive calculation, $\widehat{\Psi}^{[1]}_{3|1}$ is already available. To calculate $\widehat{\Psi}^{[1]}_{5|3}$ using (7), define the artificial state $0^{*}=\{0,1,2,3,4\}$. Moreover, for the non-trivial case, the fractional weight $\widehat{\phi}_{i3}$ is obtained by estimating the probability of transitioning from another artificial state defined by pooling states $\{0,1,2\}$, to state 3.
0. Healthy 1. Mild COPD 2. Death without COPD 3. Moderate COPD 4. Death following mild COPD 5. Severe COPD 6. Death following moderate COPD
Figure 2: A seven-state multistate system.
4 Product limit estimators
We propose a novel estimator for estimating the conditional state occupation probability $\Psi_{k|j}$. This estimator is defined in terms of marginal state occupation probabilities. Due to the tree structure of the multistate system, any state will be reached from the initial state 0 by a unique path. Define $\mathcal{S}^{k}$ as the set of states $\ell$ such that state $k$ is on the path from 0 to $\ell$. Then, the conditional quantity, $\Psi_{k|j}$, can be expressed as the ratio of two marginal functions given by
$\Psi_{k|j}(t)=\frac{\mathbb{P}(S(t)\in\mathcal{S}^{k})}{\mathbb{P}(S(\infty)\in\mathcal{S}^{j})}$ (9)
where the numerator is the marginal occupation probability of occupying state $k$ or any of the other states that come after $k$ in the progressive system at time $t$, while the denominator is the marginal occupation probability of occupying state $j$ or any of the other states that come after $j$ in the system during an individual’s lifetime. We denote the resulting estimator defined by the ratio of marginal estimates of state occupation probabilities by $\widehat{\Psi}^{[2]}_{k|j}$. Since Equation (9) is a function of marginal state occupation probabilities, we begin with estimating these marginal probabilities by calculating the counting process and at-risk set using the method of Datta and Sundaram (2006), described in Section 2.1. Afterward, we obtain the Aalen-Johansen estimates of the marginal state occupation probabilities (see details below). Finally, $\widehat{\Psi}^{[2]}_{k|j}$ is obtained by the plug-in method.
The product-limit estimator of the marginal state occupation probability relies on the Nelson-Aalen estimate of the cumulative transition intensity, denoted as $\boldsymbol{A}(t)$. Let $\widehat{\boldsymbol{A}}(t)=\{\widehat{A}_{jj^{\prime}}(t)\}$ denote the corresponding estimator given by the expression:
$\widehat{A}_{jj^{\prime}}(t)=\begin{dcases}\int^{t}_{0}\mathcal{I}\Big{(}\widehat{Y}_{j}(u)>0\Big{)}\widehat{Y}_{j}(u)^{-1}\text{d}\widehat{N}_{jj^{\prime}}(u)&j\neq j^{\prime},\\ -\sum_{j^{\prime}\neq j}\widehat{A}_{jj^{\prime}}(t)&j=j^{\prime},\end{dcases}$ (10)
where $\widehat{N}_{jj^{\prime}}(t)$ and $\widehat{Y}_{j}(t)$ respectively denote the estimators of the counting process and the at-risk set (ref: Eqns (1)-(2)). The product limit of $\widehat{\boldsymbol{A}}(t)$ leads to the (marginal) Aalen-Johansen estimator of the transition probability matrix ${\boldsymbol{P}}(s,t)$,
$\widehat{\boldsymbol{P}}(s,t)=\prod_{(s,t]}\{\boldsymbol{I}+\text{d}\widehat{\boldsymbol{A}}(u)\},$ (11)
where $\boldsymbol{I}$ denotes the $Q\times Q$ identity matrix. Consequently, the marginal estimator of the state occupation probability is given by
$\widehat{\pi}_{j}(t)=\sum^{Q}_{j^{\prime}=1}\widehat{\pi}_{j^{\prime}}(0)\widehat{P}_{j^{\prime}j}(0,t),$ (12)
where $\widehat{\pi}_{j^{\prime}}(0)=\{\widehat{Y}_{j^{\prime}}(0+)\}/\{\sum^{Q}_{j=1}\widehat{Y}_{j}(0+)\}$, and $\widehat{P}_{j^{\prime}j}(0,t)$ is the $j^{\prime}j$th element of $\widehat{\boldsymbol{P}}(0,t)$. Datta and Satten (2001) demonstrated that Equation (12) provides a consistent estimator for the state occupation probability, even in non-Markov models. This finding allows us to estimate the conditional state occupation probability without the Markov assumption using a product-limit approach similar to the Aalen-Johansen calculation.
Once we construct $\widehat{\Psi}^{[2]}_{k|j}$ using the product-limit approach, directly estimating the conditional entry distribution $F_{k|j}(t)$ becomes straightforward by substituting $\widehat{\Psi}^{[2]}_{k|j}$ into Equation (8).
4.1 A toy example
We demonstrate the formulation of the new estimator using an illustrative multistate system (refer to Figure 4) that can describe the progression of Chronic Obstructive Pulmonary Disease (COPD). This system comprises seven distinct states representing various health conditions individuals may experience over time. Commencing from the healthy state (alive and event-free), which mirrors normal lung function, individuals might either die without any COPD events or transit to the mild COPD state upon the onset of mild airflow limitation typical in early-stage COPD. Subsequently, from the mild COPD state, individuals could either decease or progress to moderate COPD as their symptoms exacerbate, and airflow limitation intensifies. Advancing from the moderate COPD state, individuals could die or further deteriorate to severe COPD, typified by notable limitations in physical activity, frequent exacerbations, and a heightened risk of respiratory failure. Our focus lies in estimating the proportion of individuals that will reach the severe COPD state given a past occupation of the mild COPD state. This conditional probability is denoted by $\Psi_{5|1}$. Additionally, we will describe the estimation of $\Psi_{3|1}$ and $\Psi_{5|3}$.
The expression in (9) leads to the respective estimators for these temporal functions:
$\widehat{\Psi}^{[2]}_{3|1}(t)$ $=\frac{\widehat{\pi}_{3^{*}}(t)}{\widehat{\pi}_{1^{*}}}=\frac{\widehat{\mathbb{P}}\Big{(}S(t)\in\{3,5,6\}\Big{)}}{\widehat{\mathbb{P}}\Big{(}S(\infty)\in\{1,3,4,5,6\}\Big{)}}$
$\widehat{\Psi}^{[2]}_{5|3}(t)$ $=\frac{\widehat{\pi}_{5^{*}}(t)}{\widehat{\pi}_{3^{*}}}=\frac{\widehat{\mathbb{P}}\Big{(}S(t)\in\{5\}\Big{)}}{\widehat{\mathbb{P}}\Big{(}S(\infty)\in\{3,5,6\}\Big{)}}$
$\widehat{\Psi}^{[2]}_{5|1}(t)$ $=\frac{\widehat{\pi}_{5^{*}}(t)}{\widehat{\pi}_{1^{*}}}=\frac{\widehat{\mathbb{P}}\Big{(}S(t)\in\{5\}\Big{)}}{\widehat{\mathbb{P}}\Big{(}S(\infty)\in\{1,3,4,5,6\}\Big{)}}.$
Let’s consider the estimators for ${\Psi}_{3|1}(t)$ and ${\Psi}_{5|1}(t)$. For $\widehat{\Psi}_{3|1}(t)$, the numerator estimates the probability of occupying the artificial state $3^{*}$ that comprises state 3 and other subsequent states along its path, at time $t$. Whereas, for $\widehat{\Psi}_{5|1}(t)$, the numerator estimates the occupation probability of state 5 at time $t$. The denominator of the respective estimators has the same form. Under the competing risk framework, the denominator simply estimates the transition probability of moving from state 0 to state 1 at some time during the observation period. In terms of state occupation probabilities, the denominator is the probability of occupying the artificial state $1^{*}$ that comprises state 1 and other subsequent states along its path at some time during the observation period.
For a simple comparison of the new approach based on the ratio of product limit estimators (PLE, hereinafter) with the approach based on the fractional risk sets (FRE, hereinafter), Figure 3 shows the curves from the estimation of the probability of occupying state 5 conditional on a previous visit of state 1, ${\Psi}_{5|1}(t)$, using simulated data. The 95% pointwise confidence interval (see Section 5) corresponding to each probability curve is also shown on the plots. Based on this simulated setting, the plots reveal the validity of both methods in estimating the true probability function. Moreso, the estimators are comparably similar. In Section 6, we carry out a thorough investigation comparing the performance of both estimators.
Figure 3: The estimates and 95% pointwise bootstrap confidence intervals for $\Psi_{5|1}$ based on simulated data from the seven-state illness death model using the two proposed methods. The dark solid line represents the true probability curve, the dark dotted line represents the estimated probability and the grey dashed lines denote the confidence intervals.
5 Bootstrap-based Confidence Intervals
The utilization of nonparametric regression in conjunction with the PAV step for formulating the estimators proposed in Section 3 and 4 imposes challenges for the asymptotic analysis for these estimators. To quantify uncertainty, a bootstrap-based procedure may provide a practical alternative. However, Faraway and Jhun (1990) pointed out that a direct application of the bootstrap will generally lead to invalid results for the “smoothed” nonparametric estimators since it cannot capture the bias component of the smoothed estimator.
[…]
$\text{Prob}(S=j)=\frac{\sum^{n}_{l=1}\mathcal{I}(S_{l}(C_{l})=j)K_{\tilde{h}}(C_{l}-C^{*}_{i})}{\sum^{n}_{l=1}K_{\tilde{h}}(C_{l}-C^{*}_{i})}.$
[…]
For $10<\alpha<$ 1, let $\widehat{\Delta}_{\alpha}(t)$ be the $(1-\alpha)$ bootstrap percentile of the distribution of
$\big{|}\sin^{-1}\big{(}\sqrt{\widehat{\theta}^{*}_{k|j}(t;h)}\big{)}-\sin^{-1}\big{(}\sqrt{\widehat{\theta}_{k|j}(t;\tilde{h})}\big{)}\big{|}$
where $\widehat{\theta}^{*}_{k|j}(t;h)$ is calculated using the bandwidth $h$ from the original estimation but is based on the bootstrap sample, while $\widehat{\theta}_{k|j}(t;\tilde{h})$ uses the new bandwidth $\tilde{h}$ but is based on the original sample. The $(1-\alpha)\times 100\%$ pointwise CI for the temporal function $\theta_{k|j}(t)$ is given by
[…]
$\quad\sin^{2}\Big{\{}\min\Big{(}\frac{\pi}{2},\sin^{-1}\Big{(}\sqrt{\widehat{\theta}_{k|j}(t;h)}\Big{)}+\widehat{\Delta}_{1-\alpha/2}\Big{)}\Big{\}}\Big{]}.$
The construction of the confidence interval described above is similar to Datta and Sundaram (2006). In the subsequent section, we conduct a Monte Carlo study to examine the behavior of the smoothed bootstrap confidence intervals.
6 Simulation study
We performed extensive simulation studies to evaluate the performance of the proposed methods based on: (i) fractional risk sets (FRE), and (ii) the ratio of product limit estimators (PLE) in calculating the conditional estimators. We considered two separate multistate systems - the five-state illness-death model shown in Figure 1, and the seven-state multistate model shown in Figure 4. For each multistate system, the exact transition times were generated from a common lognormal hazard. We considered sample sizes of 100, 200, 500, and 1000.
6.1 Five-State Illness-Death Model
Earlier, in Section 3.1, we introduced the five-state illness-death model. For our simulation experiment, our primary interest lies in estimating the probability of occupying state 3 having visited state 1, denoted by $\Psi_{3|1}$. We also study the performance of the estimators for the subdistribution function $\Psi_{3|1}(t)$, and the future entry distribution for state 3 given the prior occupation of state 1, denoted by $F_{3|1}(t)$. Our data simulation process initiates with each individual starting in state 0 at time 0, with potential transitions to either state 1 or state 2. Specifically, we simulate the scenario where 60% of individuals follow the $0\rightarrow 1$ transition path, while the remaining 40% follow the $0\rightarrow 2$ transition path. The transition out of state 0 is governed by an independent Bernoulli distribution. For those individuals transitioning to state 1, we further simulate their progression, with 60% transitioning to state 3 and 40% to state 4. This transition from state 1 is also controlled by a separate independent Bernoulli distribution. Following this simulation scheme, the true value of $\Psi_{3|1}$ is 0.6.
Let $V_{i,0}$ and $V_{i,1}$ denote the respective waiting time in states 0 and 1, for individual $i$. We simulate: $V_{i,0}\sim\text{LogNorm}(0,0.5)$ and $V_{i,1}\sim\text{LogNorm}(0,0.5)$. We study the performance of our proposed estimators under the scenarios where the random inspection time for each individual is generated from (i) the uniform distribution with a lower bound of 0 and an upper bound equal to the maximum waiting time observed across all individuals in the sample, and (ii) the Weibull distribution with (shape) $\eta=3$ and (scale) $\tau=2.5$.
6.2 Seven-State Multistate Model
Here, we consider the more complex multistate system shown in Figure 4. Primarily, we are interested in estimating the probability of occupying State 5 having visited State 1, denoted by $\Psi_{5|1}$. From Eqn 6, we have $\Psi_{5|1}=\Psi_{3|1}\times\Psi_{5|3}$. In addition to studying the performance of the estimators for $\Psi_{5|1}$, we will also study the performance of the estimators for $\Psi_{5|3}$, along with the subdistribution functions; $\Psi_{5|1}(t)$ and $\Psi_{5|3}(t)$, and the future entry distributions; $F_{5|1}(t)$, and $F_{5|3}(t)$. For the data simulation, all individuals start in state 0 at time 0, and they have two possible paths for transition: either to state 1 or state 2. We generate the data as follows: 80% of the individuals follow the $0\rightarrow 1$ path, while the remaining 20% follow the $0\rightarrow 2$ path. Among those who transitioned to state 1, there are two possible paths: 70% transition to state 3, and the remaining 30% transition to state 4. Moreover, among those that visited state 3, 60% proceeded to state 5, while the remaining 40% transitioned to state 6. Note that, the respective transitions out of state 0, state 1, and state 3 are controlled by three separate Bernoulli distributions. Following this simulation approach, the true value of $\Psi_{5|1}$ is 0.6.
To explore the performance of our proposed estimators, we simulate individual waiting times in states 0 and 1 denoted by $V_{i,0}$ and $V_{i,1}$ respectively. We simulate: $V_{i,0}\sim\text{LogNorm}(0,0.5)$, $V_{i,1}\sim\text{LogNorm}(0,0.5)$, and $V_{i,3}\sim\text{LogNorm}(0,0.7)$. We consider two scenarios where we generate the random inspection time for each individual from (i) the uniform distribution with limits 0 and the maximum waiting time across the entire sample and, (ii) the Weibull distribution with $\eta=2.5$ and $\tau=4.5$.
6.3 Simulation results
We evaluated the performance of the estimator of the conditional state occupation probability $\Psi_{k|j}$ using absolute bias. For the global assessment of the estimators of the subdistribution functions, $\Psi_{k|j}(t)$, and the future state entry time distribution, $F_{k|j}(t)$, we employed the mean absolute distance (MAD), defined as follows:
[…]
where $\widehat{\theta}$ represents the estimator based on the current status data, and $\widehat{\theta}^{\text{Emp}}$ corresponds to the empirical estimator based on complete data - this served as the benchmark. To compute the MAD, we used the empirical distribution function of the inspection times, denoted by $\theta_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{I}(C_{i}\leq x)$. When the MAD value is equal to 0, it indicates that the estimators are in complete agreement on the support of the observed inspection times. We calculated the absolute bias and MAD via Monte Carlo averaging with a replication size of 1000.
For the five-state illness death model, Table 1 shows the results for the estimators constructed with the two proposed methods under the simulation setting where the true transition times follow a lognormal hazard. The results are shown for the separate scenarios where the inspection time for each individual is generated from the uniform and Weibull distribution. For each estimator, the bias and MAD values decrease with increasing sample size, thus suggesting the consistency of the estimators. Although the performance of both methods in estimating the $\Psi_{3|1}$ is comparable in most cases, the estimator based on the FRE approach tends to have a smaller bias. This finding is consistent when we consider the other temporal functions of interest. Generally, the performance of the estimators is reasonable considering the severe censoring of the transition times.
Table 2 shows the simulation results for the seven-state multistate system where the true transition times are generated from the lognormal hazard. For each quantity of interest estimated by either the FRE or PLE approach, the performance of the estimators appears to improve with increasing sample size. Under this system, we observe that the estimators based on the FRE method slightly outperform the estimators based on the PLE method. As pointed out in Lan and Datta (2010), a possible explanation of the superiority of the FRE method over the PLE method especially for later states in the system is that the PLE-based estimation of the functions for the later states depends on the estimation done in prior states, therefore, the estimation error may propagate along the path.
Table 1: Absolute bias ($\Psi_{3|1}$) and MAD ($\Psi_{3|1}(t)$ $F_{3|1}(t)$) for the scenario where the true transition times for the five-state illness-death model were generated from the lognormal distribution. Inspection time for each individual was generated from (i) uniform and (ii) Weibull distribution. The results are shown by the methods proposed: (a) FRE, and (b) PLE. The results are based on a Monte Carlo sample size of 1000.
$\Psi_{3|1}$ $\Psi_{3|1}(t)$ $F_{3|1}(t)$
Inspection time $N$ FRE PLE FRE PLE FRE PLE
100 0.057 0.065 0.055 0.059 0.065 0.061
200 0.039 0.041 0.040 0.042 0.049 0.046
500 0.026 0.025 0.029 0.029 0.035 0.032
Uniform 1000 0.019 0.018 0.023 0.022 0.027 0.025
100 0.047 0.060 0.053 0.065 0.082 0.087
200 0.036 0.045 0.041 0.047 0.070 0.072
500 0.028 0.030 0.029 0.032 0.054 0.054
Weibull 1000 0.025 0.025 0.023 0.024 0.046 0.045
Table 2: Absolute bias ($\Psi_{5|1}\ \&\ \Psi_{5|3}$) and MAD ($\Psi_{5|1}(t),\ \Psi_{5|3}(t),\ F_{5|1}(t)\ \&\ F_{5|3}(t)$) for the scenario where the true transition times for the seven-state illness-death model were generated from the lognormal distribution. Inspection time for each individual was generated from (i) uniform and (ii) Weibull distribution. The results are shown by the methods proposed: (a) FRE, and (b) PLE. The results are based on a Monte Carlo sample size of 1000.
$\Psi_{5|1}$ $\Psi_{5|3}$ $\Psi_{5|1}(t)$ $\Psi_{5|3}(t)$ $F_{5|1}(t)$ $F_{5|3}(t)$
Inspection time $N$ FRE PLE FRE PLE FRE PLE FRE PLE FRE PLE FRE PLE
100 0.052 0.060 0.059 0.068 0.044 0.049 0.049 0.056 0.060 0.059 0.057 0.059
200 0.037 0.043 0.042 0.050 0.035 0.039 0.039 0.045 0.049 0.048 0.045 0.048
500 0.024 0.027 0.027 0.031 0.026 0.029 0.028 0.034 0.038 0.040 0.033 0.040
Uniform 1000 0.018 0.019 0.021 0.022 0.021 0.025 0.022 0.029 0.030 0.035 0.025 0.035
100 0.054 0.056 0.068 0.073 0.045 0.046 0.054 0.058 0.074 0.072 0.072 0.072
200 0.036 0.044 0.040 0.048 0.034 0.034 0.038 0.040 0.060 0.060 0.058 0.060
500 0.027 0.031 0.030 0.034 0.026 0.023 0.028 0.028 0.049 0.048 0.047 0.048
Weibull 1000 0.022 0.023 0.025 0.027 0.023 0.018 0.022 0.022 0.044 0.041 0.038 0.041
6.3.1 Coverage of smoothed bootstrap confidence intervals
For each multistate system considered in this simulation study, we evaluate the performance of the smoothed bootstrap confidence intervals. To obtain the confidence intervals, we used 1000 bootstrap resamples from the original data to calculate the bootstrap percentiles. The reported empirical coverage probabilities are based on 1000 Monte Carlo iterations at a nominal level of 0.05. Table 3 shows the empirical coverage probabilities for the estimators of $\Psi_{3|1}(t)$ based on the five-state illness death model. The coverage probabilities for the estimators of $\Psi_{5|1}(t)$ based on the seven-state illness death model are shown in Table 4. The confidence intervals for the temporal functions of interest are calculated at five fixed time points. The results are shown for the scenario where the random inspection time for each individual is generated from a uniform distribution. From Tables 3 and 4, the coverage probabilities for the estimator based on both FRE and PLE approach are reasonable.
Table 3: Coverage probabilities for the estimator of $\Psi_{3|1}(t)$ for the two proposed methods; (i) FRE, and (ii) PLE. The confidence intervals are based on 1000 bootstrap samples at a nominal level of 0.05, and the coverage results are based on 1000 Monte Carlo iterations.
FRE PLE
Sample size Sample size
Time 100 200 500 1000 100 200 500 1000
0.75 0.985 0.953 0.928 0.905 0.940 0.971 0.987 0.993
1.25 0.945 0.956 0.977 0.976 0.906 0.937 0.968 0.984
1.75 0.975 0.970 0.977 0.976 0.981 0.982 0.992 0.995
2.25 0.989 0.989 0.978 0.989 0.995 0.996 0.997 0.999
2.75 0.996 0.995 0.997 0.985 0.999 0.999 0.997 0.999
Table 4: Coverage probabilities for the estimator of $\Psi_{5|1}(t)$ for the two proposed methods; (i) FRE, and (ii) PLE. The confidence intervals are based on 1000 bootstrap samples at a nominal level of 0.05, and the coverage results are based on 1000 Monte Carlo iterations.
FRE PLE
Sample size Sample size
Time 100 200 500 1000 100 200 500 1000
2.00 0.858 0.873 0.840 0.813 0.929 0.955 0.979 0.985
3.00 0.853 0.851 0.797 0.764 0.976 0.983 0.988 0.968
4.00 0.970 0.957 0.945 0.935 0.994 0.992 0.996 0.978
5.00 0.994 0.994 0.986 0.984 0.998 0.999 0.997 0.994
6.00 0.999 1.000 0.997 0.998 0.999 0.999 0.999 1.000
7 Breast cancer example (EORTC-trial 10854)
To further illustrate the two proposed estimation methods, we re-analyzed data on 2793 breast cancer patients who participated in a trial conducted by the European Organization for Research and Treatment of Cancer (EORTC-trial 10854). The trial which involved patients with early breast cancer, aimed to compare whether surgery followed by a short intensive course of perioperative chemotherapy yields better treatment than surgery alone. A comprehensive description of the study is provided in Clahsen et al. (1996), and further findings are reported in Van der Hage et al. (2001).
Figure 4 describes the multistate system employed to illustrate our proposed methods. This system comprises nine states, including event-free and alive after surgery (state 0), loco-regional relapse following surgery (state 1), distant metastasis following surgery (state 2), simultaneous occurrence of loco-regional relapse and distinct metastasis following surgery (state 3), death without the occurrence of loco-regional relapse or distinct metastasis following surgery (state 4), distant metastasis following occurrence of loco-regional relapse after surgery (state 5), death following occurrence of loco-regional relapse after surgery (state 6), loco-regional relapse following occurrence of distant metastasis after surgery (state 7), and death following occurrence of distant metastasis after surgery (state 8). The time origin for the analysis was set as the time of surgery, meaning all individuals commenced at state 0. Transition times in the original dataset were right-censored, with follow-up times ranging from 12 to 6,756 days, and a median of 3,817 days.
To obtain current status data, the inspection time for each patient was randomly generated from a uniform distribution with a lower bound of 0 and an upper bound of 6,756 days, which is the maximum follow-up time observed across all patients in the original data. The inspection times were independent of the progression between the event states. We defined a patient’s inspection state as the state occupied at their inspection time. If the patient was right-censored, the inspection state was the most recent state they occupied before censoring. To highlight the structure of the current status data from the multistate system, we provide information for five randomly selected individuals in the supplementary material. This cross-sectional version of the data poses a severe type of censoring since we would only know the state occupied by each patient at a random inspection time rather than their transition history up to a censoring time.
0. Surgery 1. Local recurrence only 2. Distant metastasis only 3. Local recurrence and distant metastasis 4. Death without local recurrence or distant metastasis 5. Distant metastasis following local recurrence 6. Death following Local recurrence 7. Local recurrence following distant metastasis 8. Death following distant metastasis
Figure 4: Multistate system for the breast cancer study from EORTC-trial 10854.
Table 5 summarizes the distribution of the state occupation at the inspection times. As expected, most patients were inspected while occupying state 0 (alive after surgery). We are interested in estimating the probability of a patient ever reaching state 5 given a previous occupation of state 1. Inference around this quantity is crucial in breast cancer studies (Vicini et al., 2003; Murata et al., 2023) since identifying patients at higher risk of distinct metastasis following local-regional relapse informs better treatment decisions to prevent or manage metastatic disease. This conditional quantity, denoted by $\Psi_{5|1}$, was estimated using both the FRE and PLE approaches.
The estimates and corresponding 95% bootstrap confidence intervals based on the FRE and PLE approach are $0.400\ (0.241,0.570)$ and $0.433\ (0.222,0.658)$, respectively. The estimation results from both proposed methods are comparable. To provide more context to the results, we conducted a purely marginal analysis (without conditioning on a previous visit to state 1) where we estimated the probability of ever reaching state 5. The resulting estimate is $0.050$. As expected, conditioning on the previous occupation of the relapse state significantly increases the probability of ever reaching state 5.
We further estimated $\Psi_{5|1}$ using the PLE approach on the original right-censored data. This yielded an estimate of $0.344$. Notably, the estimations based on the heavily censored cross-sectional data approximate the result obtained from the right-censored data, suggesting their potential as an alternative analysis approach.
Table 5: Distribution of state occupation at inspection based on the current status data from the EORTC-trial 10854.
State 0 1 2 3 4 5 6 7 8
Frequency 1941 119 122 66 93 57 29 30 336
Percent 0.695 0.043 0.044 0.024 0.033 0.020 0.010 0.011 0.120
8 Discussion
This paper carefully employs techniques from the framework of competing risks to construct nonparametric estimators of future state entry probabilities and future state entry distributions, predicated on the prior occupation of a state, utilizing current status multistate data. The cross-sectional nature of the censored event history data brings about additional complexities when attempting to construct nonparametric estimates of temporal functions within a multistate model. We tackle this problem by harnessing ideas from nonparametric regression modeling to formulate estimators of the counting process and at-risk sets that serve as foundational elements in constructing estimators of the probability functions. To estimate the conditional quantities in the context of current status data, we employed a method (FRE) that was only available for right-censored data, and we also introduced a novel method (based on PLE). Using data from a real-world application and simulation studies for illustrative purposes, we show that both estimation methods yield valid and comparable results.
In practice, each estimation method has its benefits and challenges. Constructing the estimator based on the FRE approach could appear challenging since it requires the calculation of the fractional weights. Although the PLE method may be easier to implement once the artificial states have been specified correctly, the FRE method may be preferred for calculations involving distant states since it produces more reliable estimates for such states.
The intuitive approaches we propose provide a more direct route for analyzing the severely limited event data since they utilize the averaging principle rather than maximizing a likelihood function. Suppose we consider a less complicated setting (survival model with two states); the nonparametric maximum likelihood estimator (NPMLE) of the survival function has undesirable properties since it is undefined over a set of regions. For the general multistate model with current status data, an NPMLE approach would be more problematic.
We operate under the assumption of a progressive multistate system, structured as a directed tree, where each state is reachable via a unique path originating from the root node. This implies knowledge of path information from past states at a random inspection time, although the exact transition times remain unknown. While this assumption often holds due to the inherent structure of the multistate system, there are instances where distinctions among paths can be made by incorporating auxiliary states into an expanded multistate system. Additionally, a considerable number of inspections in a specific state are necessary to calculate the occupation probability of that state. In general, a relatively large sample size is required for valid nonparametric estimation of occupation probabilities, especially when dealing with current status data, where information is significantly limited compared to complete or right-censored data.
We could extend the proposed methods to multistate models with case-II interval-censored data by utilizing similar formulas for estimating the counting process and the at-risk sets. Since the estimators of these functionals are based on averaging, we could sum over the multiple inspection times for each individual. Ignoring the correlation among the inspection times will not inhibit the unbiasedness and consistency properties of the resulting estimators. However, the efficiency of the estimators could be improved by including a weighted variance-covariance matrix in the PAV step of the isotonic regression to account for the events recorded multiple times for each individual.
Further, researchers analyzing current status multistate data may also be interested in assessing the impact of measured risk factors on the probability of occupying a future state given a prior visit to a particular state. Regression modeling based on jackknife pseudo-values (Andersen et al., 2003) presents a flexible approach for estimating covariate effects in event history analysis. Recently, Anyaso-Samuel et al. (2023) extended the pseudo-value approach to model the effect of covariates on state occupation probabilities with multistate current status data. By jackknifing any of the estimators proposed here and utilizing the guidelines provided in Anyaso-Samuel et al. (2023), researchers can easily conduct covariate inference for the future state entry probabilities.
Data Availability Statement
The data from the breast cancer study used in Section 7 are available from the European Organization for Research and Treatment of Cancer (EORTC). Restrictions apply to the availability of these data, which were used under license in this paper.